A hydrogen atom is initially at an excite state. W
Answer #1

Energy E = hf

E = 6.626*10^-34* 1.624 *10^15

E = 1.076*10^-18 J

E = 1.076*10^-18/(1.6*10^-19)

E = 6.75 eV

no . of the Orbit n^2 = E/Eo

n^2 = 13.6/6.75

n = 1.42 = 2


radius of the Orbit n = 0.529 * n^2 /Z^2 A0

r = 0.529 * 2^2

r = 2.1 *10^-10 option C it is